let f(x)=2x+1 and g(x)=3x to the power of 2. find FoG
Answer by stanbon(48558)

let f(x)=2x+1 and g(x)=3x to the power of 2. find FoG
-------------------
FoG = F[g(x)] = f[3x] = 2(3x)+1= 6x+1
======================================
Cheers,
Stan H.